# Empirical Networks are Sparse: Enhancing Multi-Edge Models with Zero-Inflation

### v1

@@ 2.4 Zero-Inflated Degree-Corrected Stochastic Block Model (zi-DCSBM)
So far, we have presupposed that the partitioning of nodes into distinct groups is given. Nevertheless, community detection can be performed using the zero-inflated model, thereby assigning labels endogenously to the nodes based on interaction and link data. Community detection methodologies can be broadly categorised into two types: those based on quality functions [30] and those rooted in Maximum Likelihood Estimation (MLE) principles [33].
The former can be seamlessly integrated with zero-inflated models. This integration necessitates the definition of a suitable partition quality function, which can then be optimised using the parameter estimation defined herein. A classical example of this approach is modularity optimisation [30], where the partition quality function $Q$ is given by the equation:
@@
where $A_{ij}$ represents the adjacency matrix, $k_{i}$ and $k_{j}$ are the degrees of nodes $i$ and $j$, $m$ is the total number of edges, $c_{i}$ and $c_{j}$ are the communities of nodes $i$ and $j$, and $\delta$ is the Kronecker delta function. In this scenario, one would replace the expectation term $\frac{k^{\text{out}}_{i}k^{\text{in}}_{j}}{m}$ within the modularity function originating from the Chung-Lu model with that from the zi-CLCM. However, a close inspection reveals that the expected number of edges between a pair remains consistent across both models, as $q_{b_{i}b_{j}}\theta^{\text{out}}_{i}\theta^{\text{in}}_{j}\lambda_{b_{i}b_{j}}=q_{b_{i}b_{j}}\frac{k^{\text{out}}_{i}k^{\text{in}}_{j}}{mq_{b_{i}b_{j}}}=\frac{k^{\text{out}}_{i}k^{\text{in}}_{j}}{m_{b_{i}b_{j}}}$. Consequently, community detection through modularity optimisation in zero-inflated models yields analogous results to scenarios where zero-inflation is not considered.
@@ 3 Performance of Zero-Inflated Multi-Edge Models
Dataset Type $N$ $M$ $m$ $d$ $\rho$ kurtosis
HS13 High-school Contacts 327 5818 188508 0.11 3.54 1244.05
SFHH Conference Interactions 403 9565 70261 0.12 0.87 4109.62
HS12 High-school Contacts 180 2220 45047 0.14 2.80 712.33
WP Workplace Contacts 92 755 9827 0.18 2.35 880.08
WP15 Workplace Contacts 217 4274 78249 0.18 3.34 695.74
HS11 High-school Contacts 126 1709 28561 0.22 3.63 725.30
Thiers11 High-school Contacts 126 1709 28561 0.22 3.63 725.30
LyonSchool Primary School Contacts 242 8317 125773 0.29 4.31 237.41
HT09 Conference Interactions 113 2196 20818 0.35 3.29 1771.89
HO Hospital Contacts 75 1139 32424 0.41 11.68 152
KH Household Contacts 47 504 32643 0.47 30.20 38.38
BB Animal Interactions 13 78 63095 1.00 808.91 10.62
@@ 3.1 DCSBM versus zi-DCSBM
Figure 2: Percentage of sample excess kurtosis of the edge count distribution captured by DCSBM (red) and zi-DCSBM (blue) for all the Sociopatterns datasets. The expected kurtosis of each model has been computed from 100 000 realisations. Generally, zero-inflated model variants shift part of the weight of the edge count distribution to the tail, increasing the kurtosis and more closely following the empirical data.
(a) Sociopatterns HS13 Dataset
(b) Sociopatterns KH dataset
@@
Excess Kurtosis
In Fig. 2, we plot the percentage of sample excess kurtosis in the edge count distribution captured by the DCSBM (red) and zi-DCSBM (blue). The plot primarily serves as a means to compare the excess kurtosis between the two models. Generally, zero-inflated model variants allow for a shift in part of the weight of the edge count distribution from the centre towards the tail, thereby (i) increasing the kurtosis and (ii) more closely following the empirical data. In the plot, we can see that zero-inflated models mostly yield higher kurtosis where the empirical data has particularly high sample kurtosis. This indicates the presence of more extreme edge counts compared to the standard model variants, better reproducing the empirical data.
@@
These results can be confirmed quantitatively by computing the chi-squared goodness-of-fit statistics. In the case of the DCSBM, we get $1442.7$. The zi-DCSBM gives $175.91$, nearly an order of magnitude smaller than the non zero-inflated model. This supports the qualitative assessment obtained from Fig. 3(b). In the supplementary information A, we provide as supplementary figures the equivalents of Figs. 3(a) and 3(b) for the remaining 10 Sociopatterns datasets.
@@
Finally, the structure of a network has a significant impact on the dynamics running on it, influencing processes such as information diffusion and opinion formation. In denser networks, diffusion processes occur more rapidly. In the supplementary information B, we show that the DCSBM greatly overestimates the diffusion speed compared to the empirical data. Conversely, its zero-inflated counterpart consistently mitigates this issue by better preserving the structure of the empirical networks.
@@ 4 Discussion
Our study reveals significant limitations of classical multi-edge network models and the potential of zero-inflated models to overcome these challenges. We have shown that empirical multi-edge networks tend to be sparse despite having a large number of edges. This sparsity means that many edges are concentrated on a few node pairs, resulting in edge count distributions with heavy tails. These characteristics pose considerable difficulties for traditional multi-edge network models. They often fail to capture both the sparse nature and the high heterogeneity observed in real-world networks.
To mitigate these limitations, we show how classical multi-edge network models can be extended to incorporate zero-inflation. This involves introducing a mechanism that accounts for the excess number of zeroes (disconnected pairs) observed in empirical data. Zero-inflation not only helps in reproducing network sparsity but also improves the modelling of heavy-tailed edge count distributions. It achieves this by reducing the formation of many pairs with low edge counts in favour of a few pairs with high edge counts. This effectively shifts part of the weight of the edge count distribution towards larger counts. Failing to account for this ubiquitous property of real-world networks can lead to significant misrepresentation of network structures and dynamics.
@@
Our results indicate that while zero-inflation significantly enhances the fit of the models to empirical data, there remain areas for improvement. In this article, we have chosen to employ modularity maximisation to infer “model-independent” node labels. This approach allowed us to evaluate the advantages of zero-inflated models compared to their classical counterparts without the confounding effects of differing block structure. However, developing block inference algorithms tailored for zero-inflated models will yield more accurate representations of network structures. This is particularly important because the blocks derived from modularity maximisation, while useful, do not always capture the full complexity of empirical networks nor can they fully exploit the advantages of zero-inflation. Moreover, zero-inflated Poisson models are limited by the restrictive nature of the Poisson distribution. In particular, Poisson models may not adequately account for over-dispersion—a common feature of count data [37]. Employing more general distributions, such as negative binomial or generalised hypergeometric distributions, can better handle over-dispersion and improve model fit [7, 13].
@@
Our findings suggest that zero-inflated models provide a more detailed understanding of network structures, capturing both the sparsity and the extreme events reflected in the data. This aligns with the need for accurate models in various applications, such as optimising distribution systems, understanding disease spread, and analysing social behaviours. In our study, the example networks are so sparse that the necessity of zero-inflation is evident. However, in less extreme cases, determining the need for zero-inflation versus an appropriate choice of block structure becomes essential. To address this, appropriate likelihood-ratio tests and model comparison techniques have been developed for various models [7, 29, 13, 5]. These methods can help ascertain the necessity of zero-inflation by comparing the fit and performance of different models on the same dataset.
Our work bridges a critical gap in network modelling by systematically introducing zero-inflation to classical multi-edge models. This offers a more accurate framework for analysing empirical networks, thus enhancing the potential for network science to contribute to a wide range of complex systems.

### v2

@@ 2.4 Zero-Inflated Degree-Corrected Stochastic Block Model (zi-DCSBM)
Node Level Zero-Inflation
The models described by Eqs. 18 and 26 account for node heterogeneity only at the Poisson level. At the zero-inflation level, they assume a uniform process with a single parameter $q_{b}$ per block $b$ (or one single parameter for the zi-CLCM as all nodes belong to the same block). This means that the model will reproduce the expected degrees but not the number of neighbors per node. To introduce degree heterogeneity at the mixture level as well, we can modify $q_{b}$ into $q_{b}q^{\text{out}}_{i}$ $q^{\text{in}}_{j}$. The parameter estimation process follows the same procedure of the simpler models. However, now we need to optimize the likelihood equation for the $2\cdot N+B$ parameters $\boldsymbol{q}$. This increment in computational complexity (discussed in Supplementary Information D), though, is offset by a model that better reproduces the empirical data.
@@
So far, we have assumed that the partitioning of nodes into distinct groups is known. However, community detection can also be performed endogenously within the zero-inflated model, assigning labels to the nodes based on interaction and link data. Community detection methods can generally be divided into two categories: those based on quality functions [33] and those based on Maximum Likelihood Estimation (MLE) principles [36].
The former approach, based on quality functions, can be readily adapted to zero-inflated models. To do so, one require a suitable partition quality function that needs to be optimized to find optimal node partitionings. A common example is modularity optimization [33], where the partition quality function $Q$ is given by:
@@
where $A_{ij}$ represents the adjacency matrix, $k_{i}$ and $k_{j}$ are the observed degrees of nodes $i$ and $j$, $m$ is the total number of edges, $c_{i}$ and $c_{j}$ are the communities of nodes $i$ and $j$, and $\delta$ is the Kronecker delta function. The term $\frac{k^{\text{out}}_{i}k^{\text{in}}_{j}}{m}$ in $Q$ denotes the expected number of interactions between nodes $i$ and $j$ according to the a CLCM fitted to the empirical network defined by the adjacency matrix $A$ [33].
In the case of the zi-CLCM, the expected number of interactions is given by $q\theta^{\text{out}}_{i}\theta^{\text{in}}_{j}$. Substituting the MLE estimates of $\boldsymbol{\theta}$ from Eq. 25, we get that expected number of interactions according to the zi-CLCM fitted to the empirical network $A$ is also given by $\frac{k^{\text{out}}_{i}k^{\text{in}}_{j}}{m}$. Thus, the partition quality function $Q$ in Eq. 34 is equally applicable to zero-inflated models, meaning that community detection via modularity optimization is unaffected by the presence of zero-inflation.
@@ 3 Performance of Zero-Inflated Multi-Edge Models
Dataset $N$ $M$ $m$ $d$ $\rho$ kurtosis
HS13 327 5818 188508 0.11 3.54 1244.05
SFHH 403 9565 70261 0.12 0.87 4109.62
HS12 180 2220 45047 0.14 2.80 712.33
WP 92 755 9827 0.18 2.35 880.08
WP15 217 4274 78249 0.18 3.34 695.74
HS11 126 1709 28561 0.22 3.63 725.30
Thiers11 126 1709 28561 0.22 3.63 725.30
LyonSchool 242 8317 125773 0.29 4.31 237.41
HT09 113 2196 20818 0.35 3.29 1771.89
HO 75 1139 32424 0.41 11.68 152
KH 47 504 32643 0.47 30.20 38.38
BB 13 78 63095 1.00 808.91 10.62
@@ 3.1 DCSBM versus zi-DCSBM
In the Supplementary Information, we show that, despite the inherent limitations of the Poisson distribution in capturing heavy tails, introducing zero-inflation redistributes probability mass toward the tails of the edge count distribution, allowing the model to better reflect the extreme values observed in the empirical data.
@@
These results can be confirmed quantitatively by computing the chi-squared goodness-of-fit statistics. In the case of the DCSBM, we get $1442.7$. The zi-DCSBM gives $175.91$, nearly an order of magnitude smaller than the non zero-inflated model. This supports the qualitative assessment obtained from Fig. 2 (B). In the supplementary information A, we provide as supplementary figures the equivalents of Fig. 2 for the remaining 10 Sociopatterns datasets.
@@
Figure 4: The zi-DCSBM captures properties of empirical networks significantly better that its non-zero inflated variant. (top) Percentage of the empirical average clustering coefficient captured by DCSBM (red) and zi-DCSBM (blue) for all the Sociopatterns datasets. (center) Percentage of the empirical average shortest path length captured by DCSBM (red) and zi-DCSBM (blue) for all the Sociopatterns datasets. (bottom) Percentage of the empirical spectral gap captured by DCSBM (red) and zi-DCSBM (blue) for all the Sociopatterns datasets. The expected values of the properties of each model have been computed from 1 000 realisations.
Diffusion Speed
The structure of a network significantly impacts dynamics running on it, influencing processes such as information diffusion and opinion formation [4]. In denser networks, diffusion processes occur more rapidly. This can be quantified by the spectral gap–a measure of connectivity and diffusion potential–defined as the difference between the smallest and second smallest eigenvalues of the Laplacian matrix [14]. A larger spectral gap generally indicates faster diffusion across the network.
In Fig. 4 (bottom) we show the percentage of the empirical spectral gap captured by the DCSBM and its zero-inflated counterpart across all Sociopatterns datasets. Our results show that the DCSBM consistently overestimates the diffusion speed, yielding a larger spectral gap than that observed in empirical data. Conversely, the zi-DCSBM, by better preserving network sparsity, aligns more closely with the observed spectral gap values. This offers an improved representation of diffusion properties in real-world networks. In Supplementary Information B, we provide further details about this analysis.
Small-Worldness
Another key feature distinguishing empirical networks is their small-worldness: small world networks are characterized by both a high clustering coefficient and a low average path length. Small-world networks facilitate efficient communication and diffusion processes, as they maintain short paths between nodes while also forming densely connected clusters.
Due to their higher predicted density, classical network models struggle to reproduce these characteristics. In Fig. 4 (center), we show the percentage of the empirical average path length captured by the DCSBM and zi-DCSBM models across the datasets, and in Fig. 4(top), we show the percentage of the empirical clustering coefficient for both models.
Overall, our results demonstrate that the zi-DCSBM provides a substantially closer match to the empirical properties of the network analysed compared to the standard DCSBM. This improvement highlights the zi-DCSBM’s ability to better preserve network sparsity and realistic connectivity patterns, avoiding the artificial inflation of connectivity that tends to occur in classical multi-edge network models.
@@ 4 Discussion
Our study reveals significant limitations of classical multi-edge network models and the potential of zero-inflated models to overcome these challenges. We have shown that empirical multi-edge networks tend to be sparse despite having a large number of edges. This sparsity means that many edges are bundled on a few node pairs, resulting in bimodal edge count distributions with excess zeroes and heavy tails. These characteristics pose considerable difficulties for traditional multi-edge network models as they often fail to capture the sparse nature of real-world networks. Consequently, key features of empirical networks such as diffusion speed and small-worldness are misrepresented. Hence, the practical usability of these network models is limited.
To mitigate these limitations, we show how classical multi-edge network models can be extended to incorporate zero-inflation. This mechanism accounts for the excess number of zeroes (disconnected pairs) observed in empirical data. Zero-inflation not only helps reproduce network sparsity but also remedies several critical shortcomings of classical multi-edge network models: Both diffusion speed and small-worldness are better captured via zero-inflated models.
@@
Our results indicate that, while zero-inflation significantly enhances the fit of the models to empirical data, there remain areas for improvement. As shown in Supplementary Information D, one important challenge is computational complexity: the numerical maximum-likelihood estimation scales poorly with large networks, especially when considering heterogeneous zero-inflation. Moreover, in this article, we have used modularity maximization to infer “model-independent” node labels, allowing us to compare zero-inflated models with classical counterparts without the confounding effects of differing block structures. However, developing block inference algorithms specifically tailored for zero-inflated models would yield more accurate representations of network structures. This is particularly important because blocks derived from modularity maximization, while useful, do not always capture the full complexity of empirical networks and do not fully exploit the advantages of zero-inflation.
Furthermore, the zero-inflated Poisson models used here are limited by the restrictive mean-variance relationship of the Poisson distribution, which may not adequately account for over-dispersion—a common feature of count data [40]. Employing more flexible distributions, such as the negative binomial or generalized hypergeometric distributions, could improve model fit by better capturing over-dispersion, particularly in networks with more complex count structures [8, 15].
@@
Our findings suggest that zero-inflated models provide a more detailed understanding of networks, capturing both the sparsity and the extreme events reflected in the data. This aligns with the need for accurate models in various applications, such as optimising distribution systems, understanding disease spread, and analysing social behaviours. In our study, the example networks are so sparse that the necessity of zero-inflation is evident. However, in less extreme cases, determining the need for zero-inflation versus an appropriate choice of block structure becomes essential. To address this, appropriate likelihood-ratio tests and model comparison techniques have been developed for various models [8, 32, 15, 6]. These methods can help ascertain the necessity of zero-inflation by comparing the fit and performance of different models on the same dataset, providing a rigorous basis for model selection. Further work in this direction is needed to adapt these model selection techniques to high-dimensional, sparse network settings where zero-inflation may play a critical role in accurately capturing network structure.